OpenStudy (anonymous):

What are the zeros of the function x^4 -5 x^3 + 13x - 7?

OpenStudy (anonymous):

try the easy numbers first easiest is 1, but \(1-5+13-7=2\) so that is not right then try \(-1\) and get \(1+5-13-7\) bingo

OpenStudy (anonymous):

then to find the others, start by factoring \((x+1)\) out of your equation

OpenStudy (anonymous):

so is then one of the zeros is -1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh no!! i made a mistake

OpenStudy (anonymous):

oh no! lol what is it?

OpenStudy (anonymous):

\(-1\) is not a zero

OpenStudy (anonymous):

make sure you posted the correct problem the zeros are not rational, no good way to find them

OpenStudy (anonymous):

are you sure it ends with \(-7\) and not \(+7\) ?

OpenStudy (anonymous):

well the question was create your own polynomial with a degree greater than 2. Attach the graph to the word document and find the zeros of the function. I dont really know how to do it lol

OpenStudy (anonymous):

well you picked a rather lousy one you get to make up your own?

OpenStudy (anonymous):

i can show you how to pick a nice one if you like

OpenStudy (anonymous):

yes please!

OpenStudy (anonymous):

pick the zeros and work backwards for example, if you want the zeros to be \(-2, 1, 3\) then start with \[(x+2)(x-1)(x-3)\] and multiply out that will give you the zeros you want, and be easy to graph

OpenStudy (anonymous):

so i just multiply those and it will give me a function?

OpenStudy (anonymous):

yes that will give you a polynomial of degree 3 with those zeros even easier to cheat by doing this http://www.wolframalpha.com/input/?i=%28x%2B2%29%28x-1%29%28x-3%29

OpenStudy (anonymous):

that multiplies out for you you see you get \(x^3-2 x^2-5 x+6\) it also gives you a nice graph to copy (the second one)

OpenStudy (anonymous):

omg thank you so much!

OpenStudy (anonymous):

do you think you could check my work on the rest of the assignment. I think i did it right but i just want to be sure

OpenStudy (anonymous):

go ahead and post it i will look

OpenStudy (anonymous):

thank you! just a sec!

OpenStudy (anonymous):

Graph the function f(x) = (x + 3)3 by hand and describe the end behavior. (1 point) Graph the function f(x) = –x4 – 4 by hand and describe the end behavior. (1 point) Graph the function f(x) = –3x3 + 9x2 – 2x + 3 using graphing technology and describe the end behavior. (1 point) Graph the function f(x) = x4 – 7x3 + 12x2 + 4x – 12 using graphing technology and describe the end behavior. (1 point) Without using technology, describe the end behavior of f(x) = –3x38 + 7x3 – 12x + 13. (1 point) Using complete sentences, explain how to find the zeros of the function f(x) = 2x3 – 9x + 3. (2 points) Create your own polynomial with a degree greater than 2. Attach the graph to the word document and find the zeros of the function. (3 points)

OpenStudy (anonymous):

3. Graph the function f(x) = –3x3 + 9x2 – 2x + 3 using graphing technology and describe the end behavior. (1 point) • This is a cubic function. You can determine that the ends go in different directions because the degree is an odd number. The line crosses the x-axis at -1/2 and crosses the y-axis at 4. It has a positive slope. 4. Graph the function f(x) = x4 – 7x3 + 12x2 + 4x – 12 using graphing technology and describe the end behavior. (1 point) • This is a quartic function. The line crosses the x-axis at -9 and then crosses y-axis at 1. It has a positive slope. 5. This is a cubic function. It has a very steep negative slope. It crosses the y axis at 13 and the x as at 0. 6. You could graph the function and find the x intercepts or factor the equation and solve for f(x) = 0 ill post the graphs in a second

OpenStudy (anonymous):

you can check your answers by looking at wolfram first one looks just like \(y=x^3\) but shifted left 3 units

OpenStudy (anonymous):

okay, thanks!

OpenStudy (anonymous):

it even tells you to use graphing technology, so that is not cheating

OpenStudy (anonymous):

for example, the third one looks like this http://www.wolframalpha.com/input/?i=+%E2%80%933x3+%2B+9x2+%E2%80%93+2x+%2B+3

OpenStudy (anonymous):

end behaviour, starts at \(\infty\) and then goes to \(-\infty\)

OpenStudy (anonymous):

which graph is it?

OpenStudy (anonymous):

look at the top, you will see which one i graphed i think it is #3

OpenStudy (anonymous):

there is a mistake in your answer here 4. Graph the function f(x) = x4 – 7x3 + 12x2 + 4x – 12 using graphing technology and describe the end behavior. (1 point) • This is a quartic function. The line crosses the x-axis at -9 and then crosses y-axis at 1. It has a positive slope.

OpenStudy (anonymous):

it is not a line, and and therefore does not have a positive slope

OpenStudy (anonymous):

end behaviour, as \(x\to \infty\) you have \(f(x)\to \infty\) and also as \(x\to -\infty\) you have \(f(x)\to \infty\)

OpenStudy (anonymous):

similar mistake here 5. This is a cubic function. It has a very steep negative slope. It crosses the y axis at 13 and the x as at 0.

OpenStudy (anonymous):

it is a cubic function, but since it is not a line it does not have a slope

OpenStudy (anonymous):

actually, i am not sure which function #5 was referring to

OpenStudy (anonymous):

number 5 is f(x) = –3x38 + 7x3 – 12x + 13.

OpenStudy (anonymous):

it it really \(-x^{38}+7x^3-12x+13\) ???

OpenStudy (anonymous):

because that is a polynomial of degree 38, not degree 3

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

ok then change your answer the degree is 38 so it is not cubic

OpenStudy (anonymous):

so its quartic?

OpenStudy (anonymous):

because the degree is even, and the leading coefficient is negative, you know as \(x\to -\infty\) you have \(f(x)\to -\infty\) and as \(x\to \infty\) also \(f(x)\to -\infty\)

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

quartic means degree is 4

OpenStudy (anonymous):

quintic is 5

OpenStudy (anonymous):

so if its to the 38th degree what is it?

OpenStudy (anonymous):

after that you just says "nth degree" for for this one it is a polynomial of degree 38

OpenStudy (anonymous):

no special term for it

OpenStudy (anonymous):

okay so then whats the end behavior? Sorry, Im just really bad at math lol

OpenStudy (anonymous):

lol no problem end behaviour is what i wrote above, for the reason i wrote also

OpenStudy (anonymous):

so what kind of function is it again?

OpenStudy (anonymous):

lets cut to the chase if the degree is even, and the leading coefficient is positive, like \(x^4\) then the end behaviour is this: as \(x\to -\infty\) and as \(x\to \infty\) you have \(f(x)\to \infty\)

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

if the degree is even, and the leading coefficient is negative, like say \(-x^6\) then the end behaviour is this: as \(x\to \infty\) and as \(x\to -\infty\) you have \(f(x)\to -\infty\)

OpenStudy (anonymous):

that is the case for the one that started with \(-x^{38}\)

OpenStudy (anonymous):

so the end behavior for that is x→∞ and as x→−∞ you have f(x)→−∞

OpenStudy (anonymous):

now if the degree is odd, and the leading coefficient is positive, like say \(x^3\) then the end behaviour is this: as \(x\to -\infty\) you have \(f(x)\to -\infty\) and as \(x\to \infty\) you ahve \(f(x)\to \infty\)

OpenStudy (anonymous):

because its negative and even

OpenStudy (anonymous):

right

OpenStudy (anonymous):

okayyy, that makes a lot more sense

OpenStudy (anonymous):

and finally, the last case, if the degree is odd and the leading coefficient is negative, like say \(-x^7\) then as \(x\to -\infty\) you have \(f(x)\to \infty\) and as \(x\to \infty\) you have \(f(x)\to -\infty\) those are the 4 cases

OpenStudy (anonymous):